Assume that when adults with smartphones are randomly​ selected, 58% use them in meetings or classes. If 13 adult smartphone users are randomly​ selected, find the probability that fewer than 5 of them use their smartphones in meetings or classes.
Answer #1

This is a binomial distribution

p = 0.58

n = 13

P(X = x) =13Cx * 0.58x * (1 - 0.58)13-x

P(X < 5) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4)

              = 13C0 * 0.580 * 0.4213 + 13C1 * 0.581 * 0.4212 + 13C2 * 0.582 * 0.4211 + 13C3 * 0.583 * 0.4210 + 13C4 * 0.584 * 0.429

              = 0.0446